if the sum of five consecutive integers is 60. Fund the value of the largest integer.
pls. help me.

Answer by rfer(16322)

x+(x+1)+(x+2)+(x+3)+(x+4)=60
5x+10=60
5x=50
x=10
x+4=14 the largest